# The influence of the 3D Galactic gas structure on cosmic-ray transport and -ray emission

### v1

@@ 3.4 Gamma-ray emission
In Fig 14, we show the mean distributions and uncertainties of the simulated local emissivities for -decay and bremsstrahlung -ray emission, as well as for the total -ray emission (including IC). We show these results in the Galactic plane at 10āGeV. The profile of the total -ray emission is directly correlated to dense gas regions as a consequence of -decay being the dominant emission channel. Unlike the CR transport case, -decay and bremsstrahlung emission scale linearly with gas density, and the corresponding uncertainty is similar to the one of the gas reconstruction, in particular, at locations of high H2 density. Consequently, -decay and bremsstrahlung emission exhibit uncertainties of magnitudes comparable to their corresponding mean distribution.
@@
In contrast, IC emission follows from the interaction between CR electrons and the ISRF, where the latter is introduced in the simulations by a 2D model, as discussed in section 2. We show the corresponding mean distribution and uncertainty in the Galactic plane at 10āGeV and 1āTeV in Fig. 15. In this case, structures related to the 3D gas maps are faint, getting more detailed at higher energies. At 1āTeV we start to see a direct imprint of the gas distribution, which is visible at 5ākpc away from the GC in every direction. Unlike the other emission channels, IC emission does not scale linearly with the gas density. As a result, features on the emissivities only appear due to the existing structures in the distribution of CR electrons, which explains the uncertainties being negligible, as seen for the CR fluxes.
@@
In Fig 16 we compare the -ray spectra from regions of varying gas density, where we show the simulated mean total -ray together with the conrresponding uncertainties at different locations in the Galactic plane. We note that the emission is mostly proportional to the gas density, being highest at the GC. The uncertainties observed in -ray spectra are much larger than for the CR spectra and correlate directly with most regions with high gas density, where the uncertainties of the gas are also highest. The emission is however not directly correlated to the gas density everywhere. This can be the case if the source density is locally enhanced. For instance, the low gas density location at with ācm-3 features a higher total -ray emission than in the vicinity of Earth with density ācm-3. We also find that at the aforementioned low density region, the total -ray emission exhibits a change in slope when compared to the emission at the other locations. This is a consequence of the individual emission channels, which we show in Fig. 17, where we find that the slope change is related to the locally dominant IC channel. The local dominance of IC emission is a consequence of the joint effect of the ISRF being high and a locally low gas density. At Earth, -decay emission is dominant with IC emission becoming comparable in the 100-600āGeV energy range. At the GC, the gas density is so high bremsstrahlung emission dominates over IC emission. In our gas model we find a low uncertainty at the GC, which is a consequence of the underlying velocity field model and leads to negligible uncertainties of -ray emission at this exact location.
@@
In Fig 18, we show the LOS-integrated emission maps for all channels at 10āGeV and 1āTeV. We note that all major structures in the total emission follow from the dominating -decay emission. For bremsstrahlung emission, contributions are concentrated in the Galactic plane where the gas density is highest. In contrast to its corresponding local emissivity, the LOS-integrated IC emission exhibits no gas structures. Furthermore, IC emission extends to higher latitudes at 1āTeV than at 10āGeV. For reference, in Fig. 19 we show the individual HI and H2 all-sky maps. In the case of bremsstrahlung, all structures are mostly determined by the distribution of H2, as this contributes the highest density in the Galactic plane, where CR are densest. In contrast, -decay emission exhibits structures from both H2 and HI, where the high-latitude emission clearly follows the distribution of HI gas.
@@
In Fig. 20 we show -ray spectra from different locations in the sky. We particularly choose some star-forming regions and giant molecular clouds, alongside a high latitude under-dense region. Here, the locations in the sky where the total emission is highest are again proportional to the gas density, due to the dominance of -decay emission. We see that the uncertainties become negligible, as they are defined for each voxel in the 3D gas distribution and have little effect after integrating along the LOS. Furthermore, we find that the spectral slope of the -ray spectra from dense regions like Aquila, the Cepheus Flare and the under-dense region at varies slightly, due to the contribution of IC emission. We see this in detail in Fig. 21 where we show the individual -ray emission channels along the LOS from the GC, Vela and the under-dense region at . Like for the specific emissivities, at the under-dense region, the spectral slope of IC emission varies with energy and IC emission itself is comparable to -decay emission, though remaining always subdominant. We see that despite the local boost of bremsstrahlung emission at the GC, the integrated contribution from this direction does not dominate over IC emission.
@@ 4 Summary and conclusions
We studied the impact of new 3D reconstructions of Galactic HI and H2 on our understanding of the CR transport and -ray emission. For this, we prepared a 2D and a 3D model, where the 2D model corresponds to an azimuthally-averaged version of the new gas maps interpolated to the simulation grid used, while the 3D model directly interpolates the new gas maps onto the simulation grid. Then, we solved the CR transport equation using PICARD for isotopes up to silicon. Finally, we used the corresponding transport results for the computation of Galactic diffuse -ray emission. For our simulations, we used 20 possible realizations of the new gas reconstruction to determine the impact of the uncertainties of the gas distribution at all stages of the simulations.
For CR transport, we found that, to account for a 3D structured gas distribution when fitting to B/C ratio data, the diffusion coefficient and AlfvĆØn velocity needed to be decreased in comparison to the 2D model, effectively reducing the diffusive strength. Furthermore, the rigidity dependence of the diffusion coefficient needed to be adjusted.
@@
When comparing results of the axisymmetric model with the 3D model, we found a higher flux of primaries at low energies correlated to a smaller flux of secondaries at higher energies in the 3D model with respect to the 2D model. All of these effects were higher at the GC, where we noted that spectral differences come as a consequence of the distribution of gas structures, as local average gas densities remain equal.
@@
Finally, we studied -ray emission produced from the simulated CR species both locally in the Galactic plane via the LOS integrals. In the Galactic plane, we observed a direct correlation between the total -ray emission and dense gas regions. We saw a high sensitivity to gas fluctuations arising in the 3D specific emissivitiy maps of -decay and bremsstrahlung emission, as well as in the total emission. In locations of high gas density, the -decay channel always remained dominant, with bremsstrahlung emission being boosted and dominating over IC emission in such regions, like the GC. In low gas-density regions, IC became dominant, conditional to a locally high ISRF. From the LOS-integrated emission, the structure of the total emission is fully determined by the gas maps. For the individual emission channels, gas structures on IC are smoothed out.
Due to the multiple effects through which ISM gas influences the distribution of CRs and -ray emission, as well as the influence of the gas structures themselves in diffusive transport, we strongly endorse the usage of 3D structured gas maps in 3D CR transport simulations.
Acknowledgements.
@@
Data used in this study was extracted with the help of the CRDB cosmic-ray database (Maurin etĀ al., 2014, 2020, 2023). All figures in this publication have been created using matplotlibĀ (Hunter, 2007), where the healpy and HEALPix package were additionally used for some. Part of this work was supported by the German Deutsche Forschungsgemeinschaft, DFG project number 495252601. This research was funded in part, by the Austrian Science Fund (FWF) [I 5925-N]. Gordian Edenhofer acknowledges that support for this work was provided by the German Academic Scholarship Foundation in the form of a PhD scholarship (āPromotionsstipendium der Studienstiftung des Deutschen Volkesā). Philipp Frank acknowledges funding through the German Federal Ministry of Education and Research for the project ErUM-IFT: Informationsfeldtheorie für Experimente an GroĆforschungsanlagen (Fƶrderkennzeichen: 05D23EO1). Vo Hong Minh Phan acknowledges support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne UniversitĆ©.References
- Ackermann etĀ al. (2012) Ackermann, M., Ajello, M., Atwood, W.Ā B., etĀ al. 2012, ApJ, 750, 3
- Adriani etĀ al. (2022) Adriani, O., Akaike, Y., Asano, K., etĀ al. 2022, Phys.Ā Rev.Ā Lett., 129, 251103
- Adriani etĀ al. (2014) Adriani, O., Barbarino, G.Ā C., Bazilevskaya, G.Ā A., etĀ al. 2014, ApJ, 791, 93
- Aguilar etĀ al. (2023) Aguilar, M., Ali Cavasonza, L., Alpat, B., etĀ al. 2023, Phys.Ā Rev.Ā Lett., 130, 211002
- Aguilar etĀ al. (2018) Aguilar, M., Ali Cavasonza, L., Ambrosi, G., etĀ al. 2018, Phys.Ā Rev.Ā Lett., 120, 021101
- Arras etĀ al. (2019) Arras, P., Baltac, M., Ensslin, T.Ā A., etĀ al. 2019, NIFTy5: Numerical Information Field Theory v5
- Beck (2015) Beck, R. 2015, A&AĀ Rev., 24, 4
- Blasi (2013) Blasi, P. 2013, A&AĀ Rev., 21, 70
- Cerri etĀ al. (2017) Cerri, S.Ā S., Gaggero, D., Vittino, A., Evoli, C., & Grasso, D. 2017, J. Cosmology Astropart. Phys., 2017, 019
- Dampe Collaboration (2022) Dampe Collaboration. 2022, Science Bulletin, 67, 2162
- DAMPE Collaboration (2022) DAMPE Collaboration. 2022, Science Bulletin, 67, 2162
- Edenhofer etĀ al. (2024) Edenhofer, G., Frank, P., Roth, J., etĀ al. 2024, Re-Envisioning Numerical Information Field Theory (NIFTy.re): A Library for Gaussian Processes and Variational Inference
- EnĆlin etĀ al. (2009) EnĆlin, T.Ā A., Frommert, M., & Kitaura, F.Ā S. 2009, Phys.Ā Rev.Ā D, 80, 105005
- EnĆlin (2019) EnĆlin, T.Ā A. 2019, Annalen der Physik, 531, 1800127
- Evoli etĀ al. (2021) Evoli, C., Amato, E., Blasi, P., & Aloisio, R. 2021, Phys.Ā Rev.Ā D, 104, 123029
- Evoli etĀ al. (2020) Evoli, C., Blasi, P., Amato, E., & Aloisio, R. 2020, Phys. Rev. Lett., 125, 051101
- Evoli etĀ al. (2018) Evoli, C., Blasi, P., Morlino, G., & Aloisio, R. 2018, Phys.Ā Rev.Ā Lett., 121, 021102
- Evoli etĀ al. (2017) Evoli, C., Gaggero, D., Vittino, A., etĀ al. 2017, Journal of Cosmology and Astroparticle Physics, 2017, 015
- Evoli etĀ al. (2020) Evoli, C., Morlino, G., Blasi, P., & Aloisio, R. 2020, Phys.Ā Rev.Ā D, 101, 023013
- Farcy etĀ al. (2022) Farcy, M., Rosdahl, J., Dubois, Y., Blaizot, J., & Martin-Alvarez, S. 2022, MNRAS, 513, 5000
- Gabici (2022) Gabici, S. 2022, A&AĀ Rev., 30, 4
- Gabici etĀ al. (2019) Gabici, S., Evoli, C., Gaggero, D., etĀ al. 2019, International Journal of Modern Physics D, 28, 1930022
- Génolini et al. (2017) Génolini, Y., Serpico, P. D., Boudaud, M., et al. 2017, Phys. Rev. Lett., 119, 241101
- Ghelfi etĀ al. (2017) Ghelfi, A., Maurin, D., Cheminet, A., etĀ al. 2017, Advances in Space Research, 60, 833
- Giacinti & Semikoz (2023) Giacinti, G. & Semikoz, D. 2023, arXiv e-prints, arXiv:2305.10251
- Girichidis etĀ al. (2022) Girichidis, P., Pfrommer, C., Pakmor, R., & Springel, V. 2022, MNRAS, 510, 3917
- Grenier etĀ al. (2015) Grenier, I.Ā A., Black, J.Ā H., & Strong, A.Ā W. 2015, ARA&A, 53, 199
- Haverkorn etĀ al. (2019) Haverkorn, M., Boulanger, F., EnĆlin, T., etĀ al. 2019, Galaxies, 7, 17
- Heinbach & Simon (1995) Heinbach, U. & Simon, M. 1995, ApJ, 441, 209
- Hunter (2007) Hunter, J.Ā D. 2007, Computing in Science & Engineering, 9, 90
- Jacobs etĀ al. (2022) Jacobs, H., Mertsch, P., & Phan, V. H.Ā M. 2022, J. Cosmology Astropart. Phys., 2022, 024
- Jacobs etĀ al. (2023) Jacobs, H., Mertsch, P., & Phan, V. H.Ā M. 2023, MNRAS, 526, 160
- Jansson & Farrar (2012a) Jansson, R. & Farrar, G.Ā R. 2012a, ApJ, 757, 14
- Jansson & Farrar (2012b) Jansson, R. & Farrar, G.Ā R. 2012b, ApJ, 761, L11
- Jóhannesson et al. (2018) Jóhannesson, G., Porter, T. A., & Moskalenko, I. V. 2018, ApJ, 856, 45
- Kissmann (2014) Kissmann, R. 2014, Astroparticle Physics, 55, 37
- Kissmann etĀ al. (2015) Kissmann, R., Werner, M., Reimer, O., & Strong, A. 2015, Astroparticle Physics, 70, 39
- Maurin (2020) Maurin, D. 2020, Computer Physics Communications, 247, 106942
- Maurin etĀ al. (2023) Maurin, D., Ahlers, M., Dembinski, H., etĀ al. 2023, European Physical Journal C, 83, 971
- Maurin etĀ al. (2020) Maurin, D., Dembinski, H.Ā P., Gonzalez, J., MariÅ, I.Ā C., & Melot, F. 2020, Universe, 6, 102
- Maurin etĀ al. (2022) Maurin, D., Ferronato Bueno, E., & Derome, L. 2022, A&A, 667, A25
- Maurin etĀ al. (2014) Maurin, D., Melot, F., & Taillet, R. 2014, A&A, 569, A32
- Mertsch (2020) Mertsch, P. 2020, Ap&SS, 365, 135
- Mertsch, P. & Phan, V. H. M. (2023) Mertsch, P. & Phan, V. H. M. 2023, A&A, 671, A54
- Mertsch, P. & Vittino, A. (2021) Mertsch, P. & Vittino, A. 2021, A&A, 655, A64
- Moskalenko & Strong (1998) Moskalenko, I.Ā V. & Strong, A.Ā W. 1998, ApJ, 493, 694
- Moskalenko etĀ al. (2002) Moskalenko, I.Ā V., Strong, A.Ā W., Ormes, J.Ā F., & Potgieter, M.Ā S. 2002, ApJ, 565, 280
- Nakanishi & Sofue (2003) Nakanishi, H. & Sofue, Y. 2003, PASJ, 55, 191
- Nakanishi & Sofue (2006) Nakanishi, H. & Sofue, Y. 2006, PASJ, 58, 847
- Nava etĀ al. (2016) Nava, L., Gabici, S., Marcowith, A., Morlino, G., & Ptuskin, V.Ā S. 2016, MNRAS, 461, 3552
- Padovani etĀ al. (2020) Padovani, M., Ivlev, A.Ā V., Galli, D., etĀ al. 2020, SpaceĀ Sci.Ā Rev., 216, 29
- Phan etĀ al. (2023) Phan, V. H.Ā M., Recchia, S., Mertsch, P., & Gabici, S. 2023, Phys.Ā Rev.Ā D, 107, 123006
- Pohl etĀ al. (2008) Pohl, M., Englmaier, P., & Bissantz, N. 2008, ApJ, 677, 283
- Porter et al. (2022) Porter, T. A., Jóhannesson, G., & Moskalenko, I. V. 2022, The Astrophysical Journal Supplement Series, 262, 30
- Schlickeiser (2002) Schlickeiser, R. 2002, Cosmic Ray Astrophysics
- Selig etĀ al. (2013) Selig, M., Bell, M.Ā R., Junklewitz, H., etĀ al. 2013, A&A, 554, A26
- Simpson etĀ al. (2023) Simpson, C.Ā M., Pakmor, R., Pfrommer, C., Glover, S. C.Ā O., & Smith, R. 2023, MNRAS, 520, 4621
- Skilling (1975) Skilling, J. 1975, MNRAS, 173, 255
- Steininger etĀ al. (2019) Steininger, T., Dixit, J., Frank, P., etĀ al. 2019, Annalen der Physik, 531, 1800290
- Strong & Moskalenko (1998) Strong, A.Ā W. & Moskalenko, I.Ā V. 1998, ApJ, 509, 212
- Strong etĀ al. (2007a) Strong, A.Ā W., Moskalenko, I.Ā V., & Ptuskin, V.Ā S. 2007a, Annual Review of Nuclear and Particle Science, 57, 285
- Strong etĀ al. (2007b) Strong, A.Ā W., Moskalenko, I.Ā V., & Ptuskin, V.Ā S. 2007b, Annual Review of Nuclear and Particle Science, 57, 285
- Strong etĀ al. (2000) Strong, A.Ā W., Moskalenko, I.Ā V., & Reimer, O. 2000, ApJ, 537, 763
- Tan & Ng (1983) Tan, L.Ā C. & Ng, L.Ā K. 1983, Journal of Physics G: Nuclear Physics, 9, 227
- Thaler etĀ al. (2023) Thaler, J., Kissmann, R., & Reimer, O. 2023, Astroparticle Physics, 144, 102776
- Thomas & Pfrommer (2019) Thomas, T. & Pfrommer, C. 2019, MNRAS, 485, 2977
- Thomas etĀ al. (2020) Thomas, T., Pfrommer, C., & EnĆlin, T. 2020, ApJ, 890, L18
- Tibaldo etĀ al. (2021) Tibaldo, L., Gaggero, D., & Martin, P. 2021, Universe, 7, 141
- Yusifov & Küçük (2004) Yusifov, I. & Küçük, I. 2004, A&A, 422, 545

### v2

@@ 3.4 Gamma-ray emission
In Fig 11, we show the mean distributions and uncertainties of the simulated local emissivities for -decay, bremsstrahlung, and IC -ray emission, as well as for the total -ray emission (including IC). We show these results in the Galactic plane at 1 TeV. The profile of the total -ray emission is directly correlated with dense gas regions as a consequence of -decay being the dominant emission channel. -decay dominates by 2 orders of magnitude at this energy. Furthermore, -decay and bremsstrahlung emission exhibit uncertainties of magnitudes comparable to their corresponding mean distribution.
@@
In contrast to the other channels, IC emission follows from the interaction between CR electrons and the ISRF, where the latter is introduced in the simulations by a 2D model, as discussed in section 2. In this case, structures related to the 3D gas maps are faint. Unlike the other emission channels, IC emission does not scale linearly with the gas density. As a result, features on the emissivities only appear due to the existing structures in the distribution of CR electrons, which explains the uncertainties being negligible, as seen for the CR fluxes.
@@
In Fig 12 we show the -ray spectra for the total and individual -ray emission channels, where we find a different slope change at the location (,) kpc compared to the GC that is related to a local dominantion of IC emission. As expected, the emission is mostly proportional to the gas density, being highest at the GC. The uncertainties observed in -ray spectra are much larger than for the CR spectra and correlate directly to regions with high gas density, where the uncertainties of the gas are also highest. The local dominance of IC emission is a consequence of the joint effect of the ISRF being high and a locally low gas density. At Earth, -decay emission is dominant with IC emission becoming comparable in the 100-600 GeV energy range. At the GC, the gas density is so high bremsstrahlung emission dominates over IC emission. In our gas model we find a low uncertainty at the GC, which is a consequence of the underlying velocity field model and leads to negligible uncertainties of -ray emission at this exact location.
@@
In Fig 13, we show the LOS-integrated emission maps for all channels at 10 GeV and 1 TeV. We note that all major structures in the total emission follow from the dominating -decay emission. For bremsstrahlung emission, contributions are concentrated in the Galactic plane where the gas density is highest. Similar to its corresponding local emissivity, the LOS-integrated IC emission is not strongly correlated with the gas density and does not exhibit gas structures on the sky map. Furthermore, IC emission extends to higher latitudes at 1 TeV than at 10 GeV. For reference, in Fig 14 we show the individual HI and H2 all-sky maps. In the case of bremsstrahlung, all structures are mostly determined by the distribution of H2, as this contributes the highest density in the Galactic plane, where CR are densest. In contrast, -decay emission exhibits structures from both H2 and HI, where the high-latitude emission clearly follows the distribution of HI gas.
@@
In Fig 15 we show LOS-integrated spectra for the individual -ray spectra emission channels from the GC, Vela and the under-dense region on the sky at . Our predictions are compared to the Fermi-LAT PASS8 data – including point sources, unresolved sources, and the isotropic extragalactic component – extracted as in De La Torre Luque, P. et al. (2023). Like for the local emissivities in Fig 12, at the under-dense region, IC emission itself is comparable to -decay emission. We note that the uncertainties become negligible, as they are defined for each voxel in the 3D gas distribution and have little effect after integrating along the LOS. Furthermore, we find that the spectral slope of the -ray spectra from dense regions like Aquila, the Cepheus Flare and the under-dense region at varies slightly, due to the contribution of IC emission. We see that despite the local bremsstrahlung emission boost at the GC, the integrated contribution from this direction does not dominate over IC emission. In B we show the total -ray emission spectra at several additional locations, both for local emissivities and integrated along the LOS.
LOS-integrated -ray spectra can also be calculated using the CR transport results of the 2D model, using the 3D gas as a target. This approach follows earlier methods employed for modeling Galactic diffuse emission, in which two gas models are used: one to simulate CR propagation; and a different one, with an appropriate column density, is used as a target for -ray emission. In this case, deviations of up to 5% in most regions occur between 2D and 3D models, and in over-dense regions such as the vicinity of the GC such deviations can increase to 13.5%. See C for further details on how the choice of the target gas distribution influences -ray emission.
@@ 4 Summary and conclusions
We studied the impact of new 3D reconstructions of Galactic HI and H2 on our understanding of the CR transport and -ray emission. For this, we prepared a 2D and a 3D model, where the 2D model corresponds to an azimuthally averaged version of the new gas maps interpolated to the simulation grid used, while the 3D model directly interpolates the new gas maps onto the simulation grid. Then, we solved the CR transport equation using PICARD for isotopes up to silicon. Lastly, we used the corresponding transport results for the computation of Galactic diffuse -ray emission. For our simulations, we used 20 possible realizations of the new gas reconstruction to determine the impact of the uncertainties of the gas distribution at all stages of the simulations.
For CR transport, we found that, to account for a 3D structured gas distribution when fitting to B/C ratio data, the diffusion coefficient and Alfvén velocity needed to be decreased in comparison to the 2D model, effectively reducing the diffusive strength. Furthermore, the rigidity dependence of the diffusion coefficient needed to be adjusted. We found that, with a more realistic gas distribution, the adjustment of these parameters was sufficient to achieve a fit to B/C ratio data with a value of 1.47. Likewise, our predictions for antiprotons led to a value of 1.89 when comparing to data.
@@
When comparing the results of the 3D model with those of the 2D model using the same transport parameters, we found that at Earth, CR fluxes deviate noticeably at energies below 10 GeV, up to 15% for Carbon, 10% for Boron, 5% for protons, 2.5% for electrons. Such deviations doubled at the GC, where the only difference between 2D and 3D models results from the local distribution of gas, since the average gas density is identical in both cases. Hence, the local distribution of gas structures in dense gas regions appears to significantly modulate the distribution of CRs. Furthermore, species of mainly secondary nature (Boron, anti-protons, TeV electrons where secondaries dominate) exhibited deviations across all energies, being 10% in the energy range studied for antiprotons at Earth and Boron at the GC. However, the deviations observed were in general more significant for nuclei, as seen for Carbon below 10 GeV.
@@
Finally, we studied -ray emission produced from the simulated CR species both locally in the Galactic plane via the LOS integrals. In the Galactic plane, we observed a direct correlation between the total -ray emission and dense gas regions. In general, -decay and bremsstrahlung emission are highly sensitive to fluctuations in the gas density. Therefore, the uncertainties in the gas distributions used for their computation are directly translated to the computed emission, as we saw in the corresponding 3D model results. In locations of high gas density, the -decay channel always remained dominant, with bremsstrahlung emission being boosted and dominating over IC emission in such regions, like the GC. In low gas-density regions, IC became dominant, conditional to a locally high ISRF. From the LOS-integrated emission, the structure of the total emission is fully determined by the gas maps. For the individual emission channels, gas structures on IC are smoothed out.
We performed an equal-setup comparison between 3D and 2D gas maps in our models, which allowed us to quantify the contribution from the structured nature of dense gas clouds to CR fluxes and -ray emissivities. Thanks to the multiple gas realizations available from the reconstructions, we have for the first time calculated how uncertainties in such a gas model propagate to CR transport and -ray emission.
Acknowledgements
@@
Data used in this study was extracted with the help of the CRDB cosmic-ray database (Maurin et al., 2014, 2020, 2023). All figures in this publication have been created using matplotlib (Hunter, 2007), where the healpy and HEALPix package were additionally used for some. This research was funded in part, by the Austrian Science Fund (FWF) 10.55776/I5925. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. Part of this work was supported by the German Deutsche Forschungsgemeinschaft, DFG project number 495252601. Gordian Edenhofer acknowledges that support for this work was provided by the German Academic Scholarship Foundation in the form of a PhD scholarship (“Promotionsstipendium der Studienstiftung des Deutschen Volkes”). Philipp Frank acknowledges funding through the German Federal Ministry of Education and Research for the project ErUM-IFT: Informationsfeldtheorie für Experimente an Großforschungsanlagen (Förderkennzeichen: 05D23EO1). Vo Hong Minh Phan acknowledges support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne Université. The authors would like to thank Stefan Lepperdinger for his contribution in the early stages of this work.
Appendix A CR Spectra for different species
Spectra and the corresponding fractional difference between their 2D and 3D model estimations are shown in Figs 16 and 17 (Carbon and Boron), 18 (Protons), 19 (Electrons), and 20 (antiprotons).
We find flux differences up to 30% at the GC for Carbon at energies below 10 GeV/n. From Fig 1, we note that at 4 kpc from the GC there is an under-dense ring in the 2D gas distribution to which CRs can diffuse with equal probability in any direction, whereas in the 3D model the local under-dense regions are located north and south of the GC, constraining the directions in which CRs can diffuse from the GC. This same effect is barely visible for protons, up to 14% below 10 GeV/n, and is negligible for electrons, with differences up to 5% below 10 GeV/n. However, for electrons, secondaries dominate the spectrum above TeV at the GC, due to the softening of the primary injection spectrum, leading to differences increasing with energy above 10 GeV/n. Like for secondary electrons, Boron also exhibits increasing differences with energy, however, in this case the 3D model predicts fluxes up to 10% lower than the 2D model. Although Carbon and Boron are anti-correlated, their local abundances can vary differently with energy due to spallation processes, since their relevant production channels are not identical. For instance, despite 12C being the most important production channel for 11B, 16O dominates the production of Carbon via 13C (followed by 12C) which contributes weakly to the production of B isotopes (Génolini et al., 2018). Furthermore, every species incurs spallation and energy losses that are modulated by the local gas distribution, as discussed for protons in Section 3.3. Therefore, the anti-correlation between Carbon and Boron observed in our simulations is modulated by such processes and varies with energy.
For protons and electrons, the differences between models are negligible, which comes as a consequence of primaries dominating the total flux, and primary protons and electrons being used for the normalization of the fluxes. We note the discrepancy between the total electron spectrum and AMS-02 data, which can be addressed by adapting the injection spectra of electrons that were used in the simulations.
Antiprotons probe different spatial scales of the ISM than light elements such as carbon and boron, and are largely correlated to the parameter space of the propagation model (Jóhannesson et al., 2016). Being entirely produced through interactions of CR nuclei with Galactic gas, they are heavily influenced by the distribution of the latter. In Fig 20 we show the predicted antiproton spectra at Earth and at the GC. As with protons, we note that our predictions are in good agreement with available data. Here we find a value of 1.89 above 10 GeV, despite the fact that antiprotons are not used to optimize the transport parameters in this work. We note that the high gas density at the GC increases the production of tertiary antiprotons, causing them to dominate the total antiproton population at energies below GeV.
Appendix B -ray emission at different regions
In Fig 21 we compare -ray spectra from regions of varying gas density, where we show the simulated mean total -ray flux together with the corresponding uncertainties at different locations in the Galactic plane. At locations with locally low gas density, the emission is not directly correlated to the gas density, as is usually the case, because IC emission dominates over the other emission channels. For example, at where the gas density is as low as cm-3 we find a higher total -ray emission than in the vicinity of Earth with gas density cm-3. We also find that at the aforementioned low-density region, the total -ray emission exhibits a change in slope above 10 GeV when compared to the emission at the other locations, as seen before for IC emission.
In Fig 22 we show -ray spectra from different directions in the sky. We particularly choose some star-forming regions and giant molecular clouds, along with a high-latitude under-dense region. Here, the total emission is again proportional to the gas density, due to the dominance of -decay emission.
Appendix C -ray emission differences due to gas target
The computation of CR transport and -ray emission can be performed with a different gas map for each in our setup. Therefore, our model supports different options for computing the -ray emission, of which we note 3 relevant ones:
1.
CR transport and -ray emission using the 3D gas distribution (3D model)
2.
CR transport and -ray emission using the 2D gas distribution (2D model)
3.
CR transport using the 2D gas distribution and -ray emission using the 3D gas distribution (2D-3D model)
In Fig. 23 we show the corresponding sky maps for the total gas distribution in the 2D and 3D gas models. The structures seen extending to higher latitudes in the 2D model come as a consequence of gas clouds localized entirely behind the GC in the 3D gas maps contributing to densities at higher latitudes when averaging to produce the 2D model. We show the fractional differences for both the 2D and 2D-3D models versus the 3D model in Fig. 24. As expected, the 2D case deviates considerably from the 3D model, especially at dense cloud regions, at which the gamma-ray flux would be below the 3D predictions. Furthermore, the 2D-3D case varies moderately, to the same degree as the CR spectra, although it exhibits a larger deviation in the vicinity of the GC.